OpenStudy (avanti):

PRECALC QUESTION: Compute the modulus and argument of these complex numbers. a. -5i b. sqrt(10) (cos(5pi/7) +isin(5pi/7))

OpenStudy (loser66):

what is the modulus of a complex number? what is the argument of a complex number?

OpenStudy (avanti):

i think the modulus is the aboslute value. and the book only told me that the argument is theta so im not sure what to do

OpenStudy (anonymous):

@Loser66 modulus is distance from \(0\), i.e. absolute value. for example \(z=x+iy\) we have \(|z|=\sqrt{x^2+y^2}\)

OpenStudy (anonymous):

@avanti for the argument, it's the angle with respect to the \(x\)-axis, so for \(z=x+iy\) we have: